# Understanding User Mental Models in AI-Driven Code Completion Tools: Insights from an Elicitation Study

### v1

@@ 3.1 Research design
This study investigates the mental model of developers interacting with CCTs. Based on the results and limitations that emerged during the literature analysis, two key objectives guided our investigation. The first revolves around the interaction with CCTs. The second one regards the interaction with explanations provided by such tools.
To achieve this, we followed an elicitation study methodology conducted through design workshops. We have chosen this technique because, particularly when performed in workshop formats, it facilitates in-depth insights into users’ thought processes, preferences, expectations, needs, and perceptions about complex systems (similar to what has been done by [39, 40, 41]). Moreover, workshops stimulate participants to collaboratively propose, discuss, and model their ideas on code completion tools, offering rich qualitative data suited to the study’s objectives. To structure the workshop discussions, we adopted a scenario-based design approach. This solution helps participants provide their answers while referring to a realistic situation [16].
@@ 3.2 Participants
The elicitation study involved a total of 56 participants 8 female, 48 male), aged between 20 and 30 years (M = 22.88, SD = 2.63, min = 20, max = 30). They were randomly divided into 8 groups of 7 participants; this group size was chosen to enable balanced discussions where each participant could contribute their perspectives while fostering dynamic group interactions. The participants were students from both the University of Bari and the University of Salerno; 4 already had bachelor’s degrees in computer science, while the remaining were students in the third year of their bachelor’s degrees in computer science. Participants included novice (20), intermediate (32), and expert programmers (4). Their programming experience ranged from 1 year to 8 years, with familiarity in languages such as C, C++, Python, Java, C#, and JavaScript. A total of 27 participants used CCTs; 22 participants had never used CCTs, but they knew what CCTs do, while 7 participants had never used and heard about CCTs.
All relevant ethical guidelines and regulations were followed, and the study received ethical approval from the Independent Ethical Review Board of the Computer Science Department of the University of Salerno. The participants digitally provided their informed consent and could leave the study anytime.
@@ 3.4 Data Collection
The researchers analyzed the transcripts systematically using an inductive thematic analysis [42]. According to the thematic analysis protocol, the evaluators performed the following 6 steps: familiarizing with the data, generating codes, mapping codes that share similar meanings, reviewing potential themes, defining and naming themes, and producing a report. The eight guiding questions structured the analysis, with the final themes answering each question.
@@ 4.1 UI for code completion
There were differing proposals with regard to when the CCTs should provide code suggestions for participants. Three main strategies emerged. The first one consists of proactive behavior (4 groups), i.e., the CCT automatically suggests the code without users’ requests when the user is not actively typing, e.g., coding ‘breaks’ in the coding flow (“When someone is blocked, there a suggestion is given to help me” – G1). Another strategy proposed by the users regards a manual activation (5 groups) by keyboard shortcut or mouse interaction (“It must be activated manually via a shortcut” – G2). This solution leaves complete control to the user over the use of AI, which could then be less intrusive. The third solution, finally, consists of a hybrid approach (6 groups). In this solution, the participants stated that a basic behavior could involve a proactive AI strategy that intervenes only when there is a well-defined context (e.g., a long code already written) and when the suggestion is almost taken for granted and difficult to discard by the user (e.g., completion of a control structure). On the other hand, manual intervention would compensate the user for the lack of automatic AI intervention, provided that the user is sure that they require AI support (“Proactive when it’s really obvious and useful, or when the user explicitly requests it” – G8). Furthermore, several participants specify that, in general, suggestions should be generated only if there is enough coding context available (4 groups) (“It’s better not to try immediately, but take from context” – G2). This was also consistent with the perceived necessity of very precise, context-sensitive recommendations [43].
@@
The second aspect concerns how to show the suggestions. A first preference expressed by the users was for a visualization inline with distinct styling (5 groups). Specifically, code suggestions should be displayed within the main coding window, but with differences in formatting, such as unique font styles or colors (“… chosen the ‘phantom’ code suggested, it is then shown in our code, to give a ‘preview’” – G6). This method allows users to view suggestions in context while maintaining clarity, particularly when the suggested code is short (“For little code, it is convenient showing it intext” – G2). In addition, a button could be useful to switch this display on or off, i.e., to show AI and user-written code in the same or different fonts (“… turn on/off the ‘show AI/human written code’ mode…” – G8). A second proposal concerns the visualization of suggestions in a separate area from the text editor (7 groups), such as a sidebar. This modality allows users to review suggestions (especially when they are lengthy or more than one) without cluttering up the main coding space while keeping the focus on the existing code (“A separate window with a list of suggestions” – G4; “For on-demand things or lengthier suggestions, use a separate window” – G3). A third strategy involves clicking on an icon for manual activation (2 groups) and displaying it in a popup near the line where the suggestion is applied. By clicking on the icon, users can see and confirm the insertion of the suggestion, minimizing interruptions and increasing the level of control (“Maybe a ‘lightbulb’ icon that allows opening a popup” – G3). A final option concerns using chatbots (2 groups), allowing users to request suggestions in a conversational manner. The suggestion would then be entered directly into the IDE when the user accepts the chatbot’s recommendation (“It should be conversational to improve suggestions” – G3).
@@
The third aspect concerns where to display the suggestion in the user interface of the IDE. The first preference expressed by participants is to show the suggestion in a sidebar (6 groups), (“It’s better on the side, separately” – G1; “In a movable tab of the IDE” – G7). The second strategy suggested by participants is to show the suggestion in a pop-up (3 groups) near the main code (“In a togglable pop-up near the main code” – G3). Another strategy that participants came up with is to show the suggestion inline (5 groups) with the code that the developer is writing; in this case, the user could also be given a clear and immediate option to accept or reject the suggestion, e.g., an accept or reject icon (“[the suggestion] is shown within the editor, with the possibility to confirm the insertion” – G8). Finally, the last mode combines the previous ones: an adaptive visualization (2 groups), in which shorter suggestions appear inline and longer ones are presented in a separate window (“In loco or in a different area. The choice depends on the code quantity” – G2). This approach allows flexibility depending on the length and complexity of the suggestion.
The fourth aspect concerns what to show in the suggestion. Participants expressed different expectations of what constitutes valuable content in code suggestions. In general, the strategy of controlling granularity emerged (7 groups). Notably, participants expressed very contrasting needs for different levels of granularity for code suggestions: while some groups agreed on preferring suggestions up to a maximum granularity level of function (“Limit to completing the function: if it completes everything, randomness increases and precision decreases” – G1), others agreed on the usefulness of generating up to entire code files or classes (“Potentially a class” – G7). Some participants proposed inserting a slider at suggestion time to control the level of detail of the suggestions, allowing developers to switch from minimal to very long suggestions according to their needs (“Let the user choose what to suggest, for example with a slider” – G2). There was no single agreement on the minimum number of tokens that should be suggested; however, there was a certain agreement on preferring, initially, minimal suggestions with details asked on-demand (4 groups) to foster the user’s reasoning (“Only suggest the function name to insert, so that [the developer] can solve the problem on their own […] pieces of code with comments that allow us to receive cues/hints; do not suggest too much code. G6). The aspect of having different alternatives offered by the code completion tool was also brought up (3 groups) as a preferable feature to improve precision in suggestions, particularly in the case of coding challenges that the developer has never faced (“Better to have many suggestions for new things, but for more ‘precise’ things it’s better to have single suggestions” – G1). A cross-cutting aspect of the above concerns the suggestion of code accompanied by documentation (2 groups) to improve comprehension and facilitate faster decision-making about the acceptance or rejection of the suggestion (“Include references to documentation for functions” – G5). A final solution, finally, consists of suggesting code adapted to the context in which it is suggested (3 groups), including variable names or coding styles relevant to the current project. According to the participants’ proposals, the contexts could be represented by the developer’s profiles (e.g., previous projects stored in repositories) or the projects already developed by the company (“[define] the standard of the suggested code, the style (in line with my style, or with the style of my company, of other repositories)…” – G4).
@@
The last aspect investigated during the study concerned how to customize the suggestions. The first need proposed by users (6 groups) concerns personalizing the timing of suggestion activation (When) to allow each developer to tailor the responsiveness of the tool to their own workflow (“The activation timing, to disable/enable automatic hints” – G5); the need of customizing the activation timing of explanations (discussed in the next subsection) also emerged to adapt to different expertise levels of programmers (“Automatic generation of explanations is fine for novices, but could be obnoxious or even ‘outrageous’ for experts” – G2). Another aspect that emerged in this dimension concerns customization of appearance (6 groups): for example, many participants emphasized the need to customize the fonts, colors, and text size of the suggestions to ensure that they are visually distinct from the user’s code according to their preferences (“Customize character size, color, font of suggestions..” – G4). A third aspect is instead related to the granularity of suggestions (6 groups): in the What dimension, we reported the strategy of controlling granularity, which generally prescribes proposing the suggestion with different levels of granularity. Here emerged that users should be able to customize this option, e.g., the user could choose a 3, 4, 5 level scale of granularity, and for each of them, users can choose what should be suggested, e.g., the statement the user is writing, the control structure, the function, the class, or others (“Set the level and scope of suggestions (the context of action – functions, loops, operators)” – G7). The need for customization also emerged for the granularity of explanations, to adapt to the user’s knowledge by generating more or fewer comments to explain the suggested code (“… the presence of automatic comments and their quantity/frequency (every line, every function, etc.)” – G5; “Customize what is generated as explanations… Customization can be simple, such as a ‘what level of programming are you?’ initial question” – G6). Finally, the last aspect that emerged was the coding style (4 groups). In particular, participants stated that developers should adapt the tool to a certain coding style by fine-tuning it by feeding in, e.g., some of their own code repositories or their companies’ past projects as examples (“Which context to use in the code generation as a ‘background training’ … set what repositories to consider” – G2). Moreover, the developers should also be able to define quality standards to which the suggested code should adhere (“Control the quality of generated code, adherence to standards (even security ones)” – G4).

### v2

@@ 2.2 Usability studies on code completion tools
2.3 The importance of mental models
A mental model is generally a mental picture users have of how a certain system works (Norman, 1983, 2013). It includes a user’s assumptions, beliefs, and expectations while interacting with a system that will lead to their actions, interactions, and problem-solving behavior. Mental models also change over time, might contain errors, and are constrained by the user’s prior experiences or technical background. Multiple mental models of a single object may be held by one person to reflect its many functions, and two persons will not always have the same mental model. If mental models are accurate and considered in the design of a system, users can anticipate system responses, feel confident during the interaction, and adapt quickly to new features. On the other hand, gaps between the users’ mental model of the system’s functionality and its actual behavior can lead users to confusion, frustration, and errors, which will reduce their productivity and satisfaction (Doi, 2021).
CCTs supported by AI employ Machine Learning (ML) algorithms or Large Language Models and rely on probabilistic computations rather than being deterministic. For example, AI-driven suggestions may use aggregated pattern data, which are not directly readable by users. Thus, the suggestions produced might be contextually nonrelevant or impossible to interpret. If a developer’s mental model does not match the underlying mechanisms of an AI tool, users might struggle to trust the tool, interpret its suggestions accurately, or recognize when and how to use it effectively.
The need to understand mental models in the domain of AI-based CCTs is multifaceted, and different aspects can benefit from a deep understanding of users’ mental models. The first is interaction with CCTs since users might predict the tool’s behavior and know when a suggestion will be useful and when it could mislead. For instance, a tool designed considering users’ mental models might allow developers to know when the AI might make contextually relevant suggestions and when it can blunder, and so on, to keep the coding flow in hand and avoid wasting time for distraction. Another aspect that can benefit regards trust. CCTs can be largely obscure to those operating outside the normal programming logic. It has been proven that transparency and interpretability are essential for establishing user trust, and understanding the user’s mental model is essential to making users comfortable when adopting the tools (Desolda et al., 2024). This helps create trust in a tool by aligning design features with users’ mental models, making developers more likely to adopt it into their daily workflow. Another aspect that can benefit from the mental model is cognitive load: CCTs are meant to reduce repetitive tasks and cognitive load; thus, users can concentrate on more difficult problem-solving jobs. However, if users cannot intuitively understand how or why these suggestions are being made, the cognitive load can actually rise rather than fall. Refining these tools to become less disjointed in relation to user mental models means we can improve the flow of work, minimize interruptions, and ultimately help users be more productive. The last aspect is the possible support to learning: especially for novice programmers, developers can recognize common coding patterns and problem-solving techniques, enhancing both tool competency and coding proficiency. The overall improvements in systems that stem from following users’ mental models throughout the design are also a potent driver for increasing systems’ adoption and acceptability (Sergeyuk et al., 2025). A recent study by Wang et al. (2023) highlights that 54% of the developers recruited for their study are eager for better CCTs.
Exploring the users’ mental models is a complex process since these models are typically abstract, partially formed, or even subconscious. Without elicitation studies, bridging this gap is an arduous task, and they are the ones that prescribe structured methodologies to help yield, analyze, and interpret users’ cognitive models regarding system functionality. The techniques for eliciting mental models include design workshops (Jacko, 2012), interviews (Rogers et al., 2023), concept mapping (Jacko, 2012), think-aloud protocols (Lewis, 1982), and scenario-based design (Rosson and Carroll, 2012). These techniques allow researchers to design intricate representations of users’ mental models of system behavior, what they expect the system to do a priori, and how they can expect to interact with, for example, CCTs. Such insights inform tool design and can help bolster our understanding of HCI within AI-augmented coding domains towards supporting frameworks and guidelines for improved user experiences in such environments and beyond.
3 Exploratory survey of AI code completion tools
To inform the design of our elicitation study, we conducted an exploratory survey of innovative AI-assisted CCTs. Rather than aiming for comprehensive or systematic coverage, this survey aims to identify tools that are either widely adopted in the industry or that are characterized by distinctive interaction modalities, relevant to our research questions. We identified candidate CCTs through a review of recent academic publications, developer forums, official product websites, and prominent marketplace listings (e.g., Visual Studio Marketplace, GitHub repositories) as of April 2024. CCTs were included if they provided a graphical user interface and exposed features for code suggestion and/or explanation. We focused on selecting those CCTs that represented the breadth of current interface and explanation approaches encountered by developers in practice.
The principal goal was not to exhaustively catalogue the CCT landscape, but to develop a functional mapping of interaction paradigms and explanation methods. This mapping directly informed the design of our empirical study by grounding scenario construction and workshop prompts in real, contemporary CCT examples. While acknowledging limitations inherent in a non-systematic approach, this selection ensures our study addresses key facets of developer–CCT interaction as manifested in actual products.
3.1 Code completion tools analysis
The analysis of the remaining 13 CCTs providing a User Interface (UI) was guided by two main aspects: the code suggestion itself and the explanation, if any, of why that code was suggested. After familiarizing ourselves with the CCTs, we analyzed users’ interaction with them across different dimensions. To simplify the analysis and the reporting of the results, we used the 5W model by Lasswell (1974), which is adopted in various fields, such as journalism and customer analysis, and more generally in problem-solving, to analyze the complete story of a fact. A similar approach is employed in the field of HCI, where the 5W model is utilized for mental model elicitation, and comprehensive user analysis (Jacko, 2012; Desolda et al., 2017; Duric et al., 2002). Each dimension of the interaction with the tool was addressed in terms of the question of the 5W model. In this phase, the 5W model guides the analysis of the state of the art of CCTs. The questions used in this phase are similar to those presented in Section˜4.3, which are used to elicit users’ mental model: this choice allows the comparison of existing CCTs with users’ expectations. This choice allows to leverage existing tools (identified during this analysis) as design inspirations during prototype development (see Section˜7), since it allows the potential identification of tools that satisfy users’ expectations (elicited in the co-design workshops, see Section˜4). The analysis of the tools with respect to the code suggestion was done by answering these questions:
  1. 1.
    Where to show the code completion in the UI?
  2. 2.
    When should the suggestion be activated during the interaction?
  3. 3.
    How should the suggested code be shown?
  4. 4.
    What should be suggested?
  5. 5.
    How should the suggestion be customized?
  1. 6.
    Where to show the code completion explanation in the UI?
  2. 7.
    When to show the explanation during the interaction?
  3. 8.
    What should be explained?
A summary of the complete set of such models is depicted in Table˜1.
Code completion Explanation
What When How to show How to customize Where What When Where
Intellicode (Svyatkovskiy et al., 2020) lines proactive grey text curent file inline - - -
Codex (Chen et al., 2021) lines manual text chat history right side answer prompt chatbot
Copilot (Friedman, 2021) lines proactive grey text current file inline answer prompt chatbot
TravTrans (Kim et al., 2021) token proactive dropdown current file inline - - -
StarCoder (Li et al., 2023) lines proactive grey text current file inline - - -
Codegeex (Zheng et al., 2023) lines proactive grey text current file inline answer, description prompt, click chatbot
CodeWhisperer (Amazon, 2024) lines hybrid grey text current file inline source click sidebar
Codeium (Codeium, 2024) lines proactive grey text current file inline improvements, reasons click chatbot
Replit (Replit Inc., 2024) lines proactive grey text current file inline description click pointer
Cody (Sourcegraph, 2025) lines proactive grey text repository inline answer click, prompt chatbot
Tabnine (Tabnine, 2025) lines proactive grey text current file inline answer prompt chatbot
Xcode (Apple Inc., 2024) lines proactive grey text current file inline - - -
Blackbox (Blackbox AI, 2024) lines proactive grey text current file inline improvement click sidebar
3.2 Code suggestions
An important characteristic affecting the interaction with CCTs is the timing to provide suggestions (when). The analysis of the existing tools revealed two main approaches: proactive and manual. Proactive CCTs provide suggestions at a certain time without requiring a user action to trigger the completion. This type of design can be either implemented in an intrusive manner (by suggesting a completion at any time) or discreetly (by generating completions only when the context is sufficient). The manual approach, in contrast, requires the user to perform a specific action to visualize the completion. However, manual activation of CCTs is not very popular, as it is adopted in only two tools. A hybrid timing approach is also implemented, even if only in the case of CodeWhisperer, as completions are proactive but can be forced through a shortcut (Amazon, 2024).
Another aspect is the content of CCTs suggestions (what is provided as a suggestion?). Two main modalities emerged from our analysis: single token suggestions (less popular, as it was proposed only with the TravTrans prototype (Kim et al., 2021)) and multiple line completions, which are widely adopted. While the first solution is quite simple, the second one results in a feature that is hard to control and assess since the length of the suggestions may vary depending on the context and LLM characteristics. As a matter of fact, multi-line suggestion length can range from short code blocks (e.g., for loops, switches, if-else) to entire functions or classes.
The suggestion’s location (where) in the UI is also worth exploring. A predominant approach is to put completions directly inside the code (inline suggestions). The other solution instead consists of providing the suggestion in a separate space with respect to the coding area. Codex (Chen et al., 2021) is the only tool that adopts the latter modality.
It is useful to discuss how to show code completions that appear before they are accepted and integrated into the live code. A consolidated approach is to display the completed code in gray, allowing users to easily distinguish it from the live code, which is typically shown in white (or other colors, depending on the users’ preferred color scheme), before accepting the proposals of the CCT. Another method that emerged from the analysis is displaying multiple possible completions in a dropdown menu, which may be particularly convenient for single-token suggestions or very short completions, as shown by TravTrans.
Finally, it is worth investigating the customizability of CCTs, as the relevance and helpfulness of suggestions strongly depend on the type of user receiving them (how to customize). Context-aware suggestions may consider the user expertise, the specific formatting/adherence to coding standards, and the functional and non-functional requirements of specific software development scenarios to provide the most accurate code snippets possible. The strategies adopted for this aim result in three categories: personalization based on the context of the current file, personalization based on a code repository, and personalization based on chat history. The first category is the most popular, while Cody (Sourcegraph, 2025) is the only tool that managed to include a code repository for more relevant completions. Nevertheless, this last approach seems to fail for big repositories. Codex provides more aware suggestions by processing the conversation between the programmer and the chatbot as context.
3.3 Explanations of the suggested code
Some of the analyzed CCTs do not offer any explanatory features. By contrast, the remaining tools use various methods to provide developers with explanations as they write new code.
The location in the UI (where) of code explanations is an important aspect of CCTs, as it influences other interaction characteristics. Three main approaches are observed: explanations are either delivered inside a chatbot, or shown in a sidebar, or finally displayed inside the coding space. Tabnine (Tabnine, 2025), Cody (Sourcegraph, 2025), CodeGeeX (Zheng et al., 2023), Copilot (Friedman, 2021), Codeium (Codeium, 2024), and Codex (Chen et al., 2021) implement the first approach, providing chat-based explanations. In contrast, CodeWhisperer (Amazon, 2024) and Blackbox (Blackbox AI, 2024) report code explanations in a sidebar. In this case, the user is not allowed of submitting prompts for additional details, but such an approach delivers code explanations quicker, as the user does not have to think about a prompt in order to receive explanations. Finally, Replit (Replit Inc., 2024) follows a unique approach, showing explanations inside a tooltip at the mouse pointer’s location (which means the explanations appear within the coding space).
@@ 4 Methodology
This section details the methodology that was employed in this study. Leveraging the 5W model by Lasswell (1974), we conducted co-design workshops that aimed at eliciting users’ mental models and expectations for CCTs.
@@ 4.1 Research design
This study investigates the mental model of developers interacting with CCTs. Based on the results and limitations that emerged during the literature analysis, we address the following research questions:
  1. RQ1:
    How do developers mentally model their interaction with AI-driven CCTs?
  2. RQ2:
    What are developers’ expectations regarding the delivery of explanations alongside AI-generated code suggestions?
  3. RQ3:
    How do developers perceive and articulate their needs for personalization in CCTs?
  4. RQ4:
    What human-centered design principles emerge for aligning CCTs with developer mental models?
To answers these questions, we followed an elicitation study methodology conducted through design workshops. We have chosen this technique because, particularly when performed in workshop formats, it facilitates in-depth insights into users’ thought processes, preferences, expectations, needs, and perceptions about complex systems (similar to what has been done by Desolda et al. (2017); Marquardt et al. (2012); Voida et al. (2005)). Moreover, workshops stimulate participants to collaboratively propose, discuss, and model their ideas on code completion tools, offering rich qualitative data suited to the study’s objectives. To structure the workshop discussions, we adopted a scenario-based design approach. This solution helps participants provide their answers while referring to a realistic situation (Rosson and Carroll, 2012).
@@ 4.2 Participants
Figure 1: Participants’ demographic information
The elicitation study involved a total of 56 participants (8 female, 48 male), aged between 20 and 30 years ($M=22.88$, $SD=2.63$). They were randomly divided into 8 groups of 7 participants; this group size was chosen to enable balanced discussions where each participant could contribute their perspectives while fostering dynamic group interactions. Figure˜1 presents participants’ demographic information. The participants were students from both the University of Bari and the University of Salerno; 4 already had bachelor’s degrees in computer science, while the remaining were students in the third year of their bachelor’s degrees in computer science. Participants included novice (20), intermediate (32), and expert programmers (4). Their programming experience ranged from 1 year to 8 years ($M=3.36$, $SD=2.07$), with familiarity in languages such as C/C++ (most common ones, as 100% of participants were familiar with them), followed by Javascript (known by 66% of participants), Python, Java, and C#. A total of 27 participants used CCTs; 22 participants had never used CCTs, but they knew what CCTs do, while 7 participants had never used and heard about CCTs.
All relevant ethical guidelines and regulations were followed, and the study received ethical approval from the Independent Ethical Review Board of the Computer Science Department of the University of Salerno. The participants digitally provided their informed consent and could leave the study at any time.
@@ 4.4 Data collection
The researchers analyzed the transcripts systematically using an inductive thematic analysis (Braun and Clarke, 2006). According to the thematic analysis protocol, the evaluators performed the following 6 steps: familiarizing with the data, generating codes, mapping codes that share similar meanings, reviewing potential themes, defining and naming themes, and producing a report. The eight guiding questions structured the analysis, with the final themes answering each question.
@@ 5.1 UI for code completion
The results of this subsection address RQ1, since it focuses on how developers mentally model their interaction with AI-driven CCTs. The findings describe developers’ preferences on when and how code suggestions should be triggered, displayed, and customized, in order to adhere to users’ internal representations and expectations toward CCT behavior. The themes, represented in Fig.˜2, also partially cover RQ3, specifically on the developers’ needs for customizing activation and display strategies.
@@
When to show suggestions
There were differing proposals with regard to when the CCTs should provide code suggestions for participants. Three main strategies emerged:
  • Proactive (4 groups): the CCT automatically suggests the code without users’ requests when the user is not actively typing, e.g., coding ‘breaks’ in the coding flow (“When someone is blocked, there a suggestion is given to help me” – G1)
  • Manual (5 groups): activation by keyboard shortcut or mouse interaction (“It must be activated manually via a shortcut” – G2). This solution leaves complete control to the user over the use of AI, which could then be less intrusive.
@@
  • Hybrid (6 groups): a basic behavior could involve a proactive AI strategy that intervenes only when there is a well-defined context (e.g., a long code already written) and when the suggestion is almost taken for granted and difficult to discard by the user (e.g., completion of a control structure). On the other hand, manual intervention would compensate the user for the lack of automatic AI intervention, provided that the user is sure that they require AI support (“Proactive when it’s really obvious and useful, or when the user explicitly requests it” – G8). Furthermore, several participants specify that, in general, suggestions should be generated only if there is enough coding context available (4 groups) (“It’s better not to try immediately, but take from context” – G2). This was also consistent with the perceived necessity of very precise, context-sensitive recommendations (Asaduzzaman et al., 2014).
How to show suggestions
The second aspect concerns how to show the suggestions. Four suggestions emerged:
@@
  • Inline with distinct styling (5 groups): code suggestions should be displayed within the main coding window, but with differences in formatting, such as unique font styles or colors (“… chosen the ‘phantom’ code suggested, it is then shown in our code, to give a ‘preview’” – G6). This method allows users to view suggestions in context while maintaining clarity, particularly when the suggested code is short (“For little code, it is convenient showing it intext” – G2). In addition, a button could be useful to switch this display on or off, i.e., to show AI and user-written code in the same or different fonts (“… turn on/off the ‘show AI/human written code’ mode…” – G8).
  • In a separate area (7 groups): visualization of suggestions in a separate part from the text editor, such as a sidebar. This modality allows users to review suggestions (especially when they are lengthy or more than one) without cluttering up the main coding space while keeping the focus on the existing code (“A separate window with a list of suggestions” – G4; “For on-demand things or lengthier suggestions, use a separate window” – G3).
  • Clicking on an icon (2 groups): manually activate suggestions by clicking on a button and displaying it in a popup near the line where the suggestion is applied. By clicking on the icon, users can see and confirm the insertion of the suggestion, minimizing interruptions and increasing the level of control (“Maybe a ‘lightbulb’ icon that allows opening a popup” – G3).
  • Using chatbots (2 groups): allow users to request suggestions in a conversational manner. The suggestion would then be entered directly into the IDE when the user accepts the chatbot’s recommendation (“It should be conversational to improve suggestions” – G3).
Where to show suggestions
The third aspect concerns where to display the suggestion in the user interface of the IDE. Discussions with participants led to four strategies:
  • In a sidebar (6 groups): the first preference expressed by participants is to show the suggestion in a sidebar (“It’s better on the side, separately” – G1; “In a movable tab of the IDE” – G7).
  • in a pop-up (3 groups): the second strategy suggested by participants is to show the suggestion in a pop-up near the main code (“In a togglable pop-up near the main code” – G3).
  • Inline(5 groups): another strategy that participants came up with is to show the suggestion inline with the code that the developer is writing; in this case, the user could also be given a clear and immediate option to accept or reject the suggestion, e.g., an accept or reject icon (“[the suggestion] is shown within the editor, with the possibility to confirm the insertion” – G8).
@@
  • Adaptive visualization (2 groups): the last mode combines the previous ones in which shorter suggestions appear inline and longer ones are presented in a separate window (“In loco or in a different area. The choice depends on the code quantity” – G2). This approach allows flexibility depending on the length and complexity of the suggestion.
What to suggest
The fourth aspect concerns what to show in the suggestion. Participants expressed different expectations of what constitutes valuable content in code suggestions:
  • Suggestions of various granularity (7 groups): in general, the strategy of controlling granularity emerged. Notably, participants expressed very contrasting needs for different levels of granularity for code suggestions: while some groups agreed on preferring suggestions up to a maximum granularity level of function (“Limit to completing the function: if it completes everything, randomness increases and precision decreases” – G1), others agreed on the usefulness of generating up to entire code files or classes (“Potentially a class” – G7). Some participants proposed inserting a slider at suggestion time to control the level of detail of the suggestions, allowing developers to switch from minimal to very long suggestions according to their needs (“Let the user choose what to suggest, for example with a slider” – G2).
  • Minimal suggestions with details asked on-demand (4 groups): there was no single agreement on the minimum number of tokens that should be suggested; however, there was a certain agreement on preferring, initially, minimal suggestions with details asked on-demand to foster the user’s reasoning (“Only suggest the function name to insert, so that [the developer] can solve the problem on their own […] pieces of code with comments that allow us to receive cues/hints; do not suggest too much code. G6).
  • Different alternatives (3 groups): the aspect of having different alternatives offered by the code completion tool was also brought up as a preferable feature to improve precision in suggestions, particularly in the case of coding challenges that the developer has never faced (“Better to have many suggestions for new things, but for more ‘precise’ things it’s better to have single suggestions” – G1).
  • Code accompanied by documentation (2 groups): a cross-cutting aspect of the above concerns the presence of documentation together with the code to improve comprehension and facilitate faster decision-making about the acceptance or rejection of the suggestion (“Include references to documentation for functions” – G5).
  • Code adapted to the context (3 groups): code should be aware of the context in which it is suggested , including variable names or coding styles relevant to the current project. According to the participants’ proposals, the contexts could be represented by the developer’s profiles (e.g., previous projects stored in repositories) or the projects already developed by the company (“[define] the standard of the suggested code, the style (in line with my style, or with the style of my company, of other repositories)…” – G4).
How to customize suggestions
The last aspect investigated during the study concerned how to customize the suggestions. Several suggestions emerged:
  • Personalize activation timing of suggestions and explanations (6 groups): the first need proposed by users concerns personalizing when to suggest explantions to allow each developer to tailor the responsiveness of the tool to their own workflow (“The activation timing, to disable/enable automatic hints” – G5). The need of customizing when to show explanations also emerged to adapt to different expertise levels of programmers (“Automatic generation of explanations is fine for novices, but could be obnoxious or even ‘outrageous’ for experts” – G2).
  • Customize appearance (6 groups): another aspect that emerged in this dimension concerns customization of appearance: for example, many participants emphasized the need to customize the fonts, colors, and text size of the suggestions to ensure that they are visually distinct from the user’s code according to their preferences (“Customize character size, color, font of suggestions..” – G4).
@@
  • Customize granularity of suggestions and explanations (6 groups): in the What dimension, we reported the strategy of controlling granularity, which generally prescribes proposing the suggestion with different levels of granularity. Here emerged that users should be able to customize this option, e.g., the user could choose a 3, 4, 5 level scale of granularity, and for each of them, users can choose what should be suggested, e.g., the statement the user is writing, the control structure, the function, the class, or others (“Set the level and scope of suggestions (the context of action – functions, loops, operators)” – G7). The need for customization also emerged for the granularity of explanations, to adapt to the user’s knowledge by generating more or fewer comments to explain the suggested code (“… the presence of automatic comments and their quantity/frequency (every line, every function, etc.)” – G5; “Customize what is generated as explanations… Customization can be simple, such as a ‘what level of programming are you?’ initial question” – G6).
  • Tailor coding style (4 groups): participants stated that developers should adapt the tool to a certain coding style by fine-tuning it by feeding in, e.g., some of their own code repositories or their companies’ past projects as examples (“Which context to use in the code generation as a ‘background training’ … set what repositories to consider” – G2). Moreover, the developers should also be able to define quality standards to which the suggested code should adhere (“Control the quality of generated code, adherence to standards (even security ones)” – G4).